Two equally charged identical metal spheres A and B repel each other with a force F another identical uncharged sphere is touched to A and then placed midway between A nad B.The net foce on C is in the direction?????
 

Grade:12

1 Answers

Ravneet
5 years ago
THEORY
  • let two metal sphere are given charge q 
  • now an uncharged identical sphere is touched with A . after touching sphere A and C will get new charge 
  • new charge = (charge on A + charge on C ) / 2
  • new charge on A = new charge on B = q/2 
NOW, sphere C with charge q/2 is placed in midway between sphere A (with charge q/2 ) and sphere B (with charge q)
Force on C due to B = KQ2/2R2     WHERE R is the distance between B and C  (greater in magnitude )
Force on C due to A = KQ2/4R2     WHERE R is the distance between A and C 
Net force =  Repulsion Force on C due to B  –  Repulsion  Force on C due to A
after solving 
Net force = KQ2/4R2 
REMEMBER: if we calculate force between A and B when charge C is not there it will be KQ2/(2R)2 = F 
so we can say Net force = F ( toward A )
becouse Repulsion force of B on C is greater